OpenStudy (anonymous):

If f(x) = 3x - 1 and g(x) = x2 - 2x - 2, what is f(g(x))

OpenStudy (anonymous):

For x2, do you mean x^2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

For this problem all you do is insert g(x) into f(x) and you get f(g(x))=(3(x^2-2x-2)-1 unless you know the point f(g(x)) is supose to equal thats as far as you can go

OpenStudy (anonymous):

or what x is suppose to equal

OpenStudy (anonymous):

wait never mind to find x you set the two equations equal to each other and you solve

OpenStudy (anonymous):

Ok. \[f(g(x))=f(x ^{2}-2x-x)=3(x ^{2}-2x-x)-1\] Then simplify from here.